CREAM Flight 2008
CSBF Integration Status, August 1, 2008
- Details
Here is a summary of today's activites at CSBF.
- The preparation of the hang test started at 6 AM (CDT).
- The weight of the payload w/o rotator and with the bottom solar panels added was measured as 3605 lbs before moving out for the hang test.
- After having the payload on the high dolly, it is found that the lift crane is too big to lift the payload with the bottom solar panels. So the bottom solar panels were removed and re-assembled at the outside for the hang test.
- The hang test in solar panel power mode was started around 9 AM (CDT).
- At the beginning of the hang test, the payload related function was tested with the instrument was in flight mode with the muon trigger.
After the function test, the instrument LPT was completed except the shutdown safe of the SFC1 at the end.
CSBF Integration Status, July 25, 2008
- Details
Here is a summary of today's activites at CSBF.
- There was no power stoppage during the overnight muon run again.
- Test data in 24th and 25th were transferred to the cdps3.
Debugging of the reduced hsk data continued today by Prashant with the Iridium connection from the CDM. Finally, he came up with a solution with the problem. Now the reduced hsk in the Iridium mode works even with two different sets, one original and the other for the CER and TCD. The details will be explained by Prashant in a separate e-mail.
CSBF Integration Status, July 24, 2008
- Details
Dear colleagues,
Here is a summary of today's activites at CSBF.
- There was no power stoppage during the overnight muon run again.
- The CPT with CDM, which consists of 4 power modes, completed tonight.
- The remaining LPT of sub-detectors for the first CPT (GSE power supply) started from morning and ended at about 2 PM with an updated manual.
During the LPT, the data flow to the UMD SOC started but stopped around noon due to a rain outside at CSBF. No more connection until the end of the CPT. JiHye got more experience with this test especially on the data relay server setup at the cdps3. Plan to have an SOC expert training next week. Candidates are Dr. Kim, Opher, Ki-Chun and Brian. Other undergraduates can join the training also.
CSBF Integration Status, July 22, 2008
- Details
Here is a summary of today's activites at CSBF.
- There was no power stoppage during the overnight muon run again. Larry told me that the heat blocking by the card board could reset the power supply. Also suggested to have the UPS connected to the GSE notebook computer so that the UPS power can be monitored. Or the Power GSE Labview program needs to have logging of the power data on. I will try to do both.
- Chris Field (CSBF) confirmed that CSBF power monitoring does not show any power off/on at that time. There are many CDM equipment with power on for 24 hours with power outlet next to what we are using and they did not have the problem.
CSBF Integration Status, July 21, 2008
- Details
Here is a summary of today's activites at CSBF.
- On Sunday early morning around 6:30 AM (CDT), there was a kind of power stoppage for about guesstimated 10-15 minutes and the muon run was stopped. The SFC was rebooted and cdaq_serv program was started but not the instrument power and HV which requires commanding. The data were copied to cdps3.
- MXR box cabling and labeling were done.



